1. A flywheel, in the form of uniform disk, is 4.0 m in diameter and

weighs 637 N. What will be its angular acceleration if it is

activated upon by net torque of 225J?



2. When you switch your room fan from medium to high speed, the

blades accelerate at 1.2 radians per second squared for 1.5

seconds. If the initial angular speed of the fan blades is 3.0

radians per second, what is the final angular speed of the fan

blades in radians per second?


1
Expert's answer
2022-01-12T08:29:55-0500

1) The angular acceleration is


"\alpha=\dfrac\tau I=\dfrac{8\tau}{md^2}=1.73\text{ rad/s}^2."

2) The final angular speed is


"\omega_f=\omega_i+\alpha t=4.8\text{ rad/s}."